OpenStudy (anonymous):

How do I know the porcentage of a fraction And how do I turn itinot a decimal?

OpenStudy (mathstudent55):

To turn a fraction into a percentage, set the fraction equal to a ratio out of 100. Then solve the proportion. For example, what percent is 5/8? Set 5/8 equal to a fraction with denominator 100 and unknown numerator, x, or x/100. Now set the fraction equal to the ratio out of 100 and solve for x: 5/8 = x/100 8x = 5 x 100 8x = 500 x = 62.5, so 5/8 = 62.5%

OpenStudy (mathstudent55):

To write a fraction as a decimal, just do long division: 5/8 = 5 divided by 8, or 8 into 5. Do that long division, and you'll get the answer